(a) Find the volume of water in a swimming pool with vertical ends and sides. The length measured at the water line is 50 ft. and the breadth is 20 ft. The bottom of the swimming pool is a plane slopping gradually downward so that the depth of the water at one end is 8 ft. and at the other end is 4 ft.


(b) If the sides, ends, and bottom of the swimming pool are constructed of tile blocks whose glazed surface dimensions are 3 in. by 6 in., and if the ends and sides of the pool extend 2 ft. above the water level, find the number of blocks used if 1/20 of the surface area is covered by sealing material.



1
Expert's answer
2021-12-06T09:45:28-0500

a)


"V=0.5(50)(20)(8+4)=6000\ ft^3"


b)


"(3)(6)\frac{1}{12^2}N\\=0.95((50)(10+6)+(20)(10+6)+(50)(20))\\N=16112"